### v2

# Why Isn’t Relational Learning Taking Over the World?

Abstract
Artificial intelligence seems to be taking over the world with systems that model pixels, words, and phonemes. The world is arguably made up, not of pixels, words, and phonemes but of entities (objects, things, including events) with properties and relations among them. Surely we should model these, not the perception or description of them. You might suspect that concentrating on modeling words and pixels is because all of the (valuable) data in the world is in terms of text and images. If you look into almost any company you will find their most valuable data is in spreadsheets, databases and other relational formats. These are not the form that are studied in introductory machine learning, but are full of product numbers, student numbers, transaction numbers and other identifiers that can’t be interpreted naively as numbers. The field that studies this sort of data has various names including relational learning, statistical relational AI, and many others. This paper explains why relational learning is not taking over the world – except in a few cases with restricted relations – and what needs to be done to bring it to it’s rightful prominence.
Introduction
AI has hit the news recently with models that learn to predict images, text, sound and video. Many have even speculated that such technologies will lead to artificial general intelligence (AGI), and be more intelligent than people.
In my AI courses I often include the clicker question of Figure 1. Reading machine learning books and papers would lead one to assume the answer is A. Reading the press on generative AI would lead one to think the answer is B. Very few of the students choose A or B.
What is the real world made of?
  • A
    Features or random variables
  • B
    Words, pixels, phonemes …
  • C
    Entities and events (e.g., plants, people, diseases, lectures, university course)
  • D
    Huh? There is a real world?
Figure 1: A clicker question
Learning and reasoning about entities, events and relations between them is central to AI:
The mind is a neural computer, fitted by natural selection with combinatorial algorithms for causal and probabilistic reasoning about plants, animals, objects, and people.
– Steven Pinker (1997)
Generative AI is characterized by models that learn from huge datasets to predict probabilities from prompts and sampling from the distributions. For example, large language models can be seen as sampling from
$P(\mbox{\emph{Next, acc}}\mid\mbox{\emph{Con}})=P(\mbox{\emph{acc}}\mid\mbox{\emph{Next, Con}})P(\mbox{\emph{Next}}\mid\mbox{\emph{Con}})$
[…]
By relational learning, I mean learning models that make probabilistic predictions about entities (things, objects, including events), their properties and relations among them. The aim is to model entities rather than modelling their manifestations in language or images.
Contemporary relational learning tools are typically trained to predict probabilities, optimizing log loss. Evaluation is much more diverse; see below.
Part of the theme of this paper is that we should not lose connection to the downstream task that a learned model is used for. Learning is usually not an end in itself, but is used for decision making. This could be for autonomous decision making, where probabilities and utilities are needed, or human-supported decision making where multiple scenarios can be evaluated by a human decision maker.
Relational Data
If you were to examine company data, environmental data, or government data – with few exceptions – you will find that the most valuable data not in natural language text or images but in spreadsheets, relational databases and other relational formats. This relational data is typically not the sort of tabular data used in introductory machine learning courses – for which standard algorithms such as gradient tree boosting (Chen and Guestrin 2016) work directly on – but is full of product numbers, student numbers, transaction numbers and other identifiers. These identifiers are often represented as numbers, but the numbers themselves don’t convey meaning. For gradient tree boosting to be effective for such data, it needs to exploit relational structure (Natarajan et al. 2012).
The simplest form of relational data is exemplified by collaborative filtering datasets such as used in the Netflix challenge (Bell and Koren 2007) and Movielens datasets (Harper and Konstan 2015). Movielens contains rating datasets containing user-movie-rating-timestamp tuples, where the user and movie are identifiers represented by arbitrary numbers. There are also tables for user properties and movie properties, including links into the Internet Movie Database (IMDB). An even simpler version is a rated dataset that ignores the actual ratings, and is just a user-movie dataset (perhaps with timestamps). To predict rating (or rated) from just the rating dataset, you can define latent properties of users and of movies and use these to predict ratings. Representing latent properties as a vector, matrix factorization (Koren, Bell, and Volinsky 2009) learns about movies and users, without learning general knowledge. Another possible technique is to do a self-join, which can result in a user-user table or a movie-movie table.
Most real-world databases are much more complicated, with multiple relations.
When considering relational data, the following questions are often implicit and affect what is represented:
  • Does the system know what exists? Most databases assume that they know what exists; things that exist are given identifiers, and what doesn’t exist is ignored. In many real-world cases, we don’t know what exists. What there a person in the house last night? Who threw the rock that broke the window? (It might have not been a rock or maybe no one threw it.)
  • Do unique identifiers denote different entities? This is known as the unique names assumption and features in logic programming and most databases. Identifying equality (two descriptions describing the same entity) arises in record linkage (Fellegi and Sunter 1969) and citation matching (Pasula et al. 2003). It is a big problem in medicine where drug dispensers need to decide whether someone asking for drugs is the same person who asked previously.
Knowledge Graphs
(Subject, verb, object) triples are the basis for knowledge graphs where the subject is an entity, the verb is a relation or property and the object is an entity or a datatype value, such as a number or a string. A collection of triples can be treated as a labelled directed graph where the entities and datatype values are nodes. There is an arc from the subject to the object, labelled with the verb. Triples are universal representations of relations: any relation can be represented as subject-verb-object triples (See e.g., Poole and Mackworth 2023, Section 16.1).
One large public knowledge graph is Wikidata (Vrandečić and Krötzsch 2014). Figure 2 (a) shows some tables representing the football (soccer) player Christine Sinclair. She is the player (man or woman) with the most goals in international play. Figure 2 (a) shows the corresponding Wikidata knowledge graph. In Wikidata, Christine Sinclair is represented by the identifier Q262802. The country Canada is represented by the identifier Q16.
entity name language
Q262802 “Christine Sinclair” en
Q16 “Canada” en
Q30 “United States of America” en
Q30 “États-Unis” fr
Q1446672 “Portland Ferns” en
member team start_time #matches #goals
Q262802 Q499946 2000 190 319
Q262802 Q1446672 2013 161 65
Figure 2: Facts about football player Christine Sinclair (Q262802): (a) in tabular form and (b) part of the corresponding Wikidata knowledge graph; from Figure 16.1 of Poole and Mackworth (2023)
This shows two ways of converting relations to triples. The first is to group columns such as “Canada”-en, giving (Q16, name, ”Canada”-en). Grouping columns does not work in general as there are exponentially many grouping of columns. It is only done in restricted cases.
The second way to converting relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). In Figure 2 (b), Q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
Another way to convert relations into triples is to pair attributes, called the star-to-clique method (Wen et al. 2016). This, however, loses information. Figure 3 shows some examples of the sort of triples that arise from this method.
Datasets
One of the most used knowledge graphs in the literature is FB15k (Bordes et al. 2013b), a large subset of Freebase. Freebase (Bollacker et al. 2008) was a free and open knowledge base developed from 2007-2016. It was acquired by Google, who donated the contents to Wikidata, and use another branch as the Google knowledge base. In building FB15k, Bordes et al. (2013b) selected entities and relations that appeared in more than 100 tuples. They also created a larger knowledge graph, FB1M, with the most frequently occurring 1 million entities, which has been used much less in subsequent research.
(A.S. Livorno Calcio, /soccer/football_team/current_roster,         Forward (association football))(Forward (association football), /soccer/football_roster_position/team,         Cambridge United F.C.)(California, /location/religion_percentage/religion, Methodism)(Ambient music, /music/genre/artists, Portishead (band))(Marriage, /people/marriage/spouse, Noel Gallagher)(Hannah Montana: The Movie,      /film/film_regional_release_date/film_release_region, Egypt)
Figure 3: Some of the test triples in FB15k. The subject and object entities have their identifier replaced with the name. The verb (property) names have been abbreviated.
Figure 3 shows some of the triples in FB15k. The last two are obviously projections of relations that are more interesting, telling us who Noel Gallagher was married to (and probably when) and when Hannah Montana: The Movie was released in Egypt.
We [^2] Bahare Fatemi, Mehran Kazemi, Ali Mehr and Martin Wang. See https://cs.ubc.ca/˜poole/res/FB15k-triples.txt for example triples for all 1344 predicates. created English version of the test set that, together with the plausible examples that existing methods ranked highly, could be used to crowdsource ground truth. The ground truth could then be used to evaluate prediction methods. However, we decided determining whether a tuple was true wasn’t a skill people could do quickly and accurately, even when they had access to Wikipedia, Google and other online sources. Positive examples were straightforward, but determining something is false requires guessing.
Toutanova and Chen (2015) noticed that a predictor could do reasonably well on FB15k by just determining that two predicates were (approximately) equivalent or inverses. To determine whether a predictor was doing more than this, they created FB15k-237 with the inverse functions removed (it would remove one first two triples of Figure 3). FB15k-237 isn’t better than FB15k; they test different things. Together they provide interesting evidence: the difference tells us whether a predictor can recognize inverses; the performance on FB15k-237 tells us whether the method can exploit other regularities. We would want a predictor that can determine inverses, but would like it to do more than that.
Including only most frequent entities means that there are no reified entities in the resulting triples. In Wikidata, around 99% of the entities appear as the subject of fewer than 10 triples (see Appendix A). Training and evaluating on datasets where all such entities are removed will not lead to better predictors on more realistic datasets.
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which not much is known. As more detail is added about an entity, there are more associated reified entities, about which there are inherently few tuples. See Appendix A.
Another standard dataset is WN18 (Bordes et al. 2013a), based on WordNet (Miller 1995), a large lexical database of English. The entities are words, and the relations include hyponym (subclass), hypernym (superclass), part_of and type, with loosely connected hierarchies for nouns, verbs, adjectives and adverbs. Most of the relations are transitive. I would not expect that words are prototypical entities.
Similar problems with standard benchmarks arise in the graph learning community (Bechler-Speicher et al. 2025; Roy et al. 2025). Although relational learning can be seen as an instance of graph learning, and graph learning an instance of relational learning, the challenges in this paper do not seem to have much in common with generic graph tasks.
JF17K (Wen et al. 2016), FB-AUTO, M-FB15K (Fatemi et al. 2020) and RelBench (Robinson et al. 2024) are benchmarks that do not convert to triples. Such datasets are a start in building more realistic test sets.
Relational predictions
As far as relational learning is concerned, it is important to distinguish two types of relational datasets.
In the complete knowledge or closed world assumption, statements not in a database are false. E.g., Airlines know the actual departure and arrival times of flights in the past, who is booked on a flight, and who showed up for boarding. They do not know who will show up tomorrow. When data is considered complete up to some time, the predictive problem is to predict the future from the past. Consider, for example, the problem of predicting the passenger load or whether I will miss a connection, of a trip from Vancouver BC to Boston MA on July 2 2026. This is a tricky prediction as July 2 is a special day – is one day after a holiday in Canada, and 2 days before one in USA – and 2026 is different in that many Canadians are boycotting USA and the FIFA World Cup is being held then; Vancouver and Boston are host cities. Rather than guessing, we should be able to make an informed prediction based on data.
Surprisingly, there is very little work on making predictions about the future from arbitrary complete databases. We can’t add an airline reservation system to a learner and get it to predict future demand. It might be because most company databases are proprietary and include much metadata, such as the database schema. Individual companies such as Amazon and Netflix have huge datasets they are mining, but the techniques are proprietary and are undoubtedly tuned to the particular task, such as maximizing profit and engagement, rather than allowing general queries of what is predicted. Steck et al. (2021) describe techniques used by Netflix. Feature engineering (such as carefully chosen joins and projections) can created tables that methods such as XGBoost work on.
The alternative to the complete knowledge assumption is the open-world assumption: the data is not complete and there is missing data. Wikipedia contains a very small proportion of the true statements even about those entities chosen to be notable enough to be included. Building predictive models from such models has two main challenges:
  • There are no negative statements.
  • Missing data is not missing at random.
Databases do not store arbitrary relations. Some relations are stored and others are derived. Which relations to store is carefully designed (Codd 1970). Database designers choose to store only true statements for some of the possible relations, selected to make the database small. They don’t represent “did-not-buy” or “has-watched-movie-starring” as explicit relations, instead choose smaller relations to explicitly represent, and use queries and rules to infer other ones.
The lack of negative statements is a feature of databases, but with the open world assumption, it makes predicting probabilities problematic. There is no way to estimate probabilities without external information about the entities and relations not represented (e.g., knowing that people are typically married to zero or one person, or that celebrities are richer than average). This meta information cannot be assumed to be explicit in a relational database.
Predicting property values, when the set of values is a constant set, such as the positions of football players, may be amenable to techniques for tabular learning, such as gradient tree boosting, following feature engineering. However, when making predictions where the values are entities (link prediction) this is infeasible.
Training
A standard to make probabilistic predictions is to use model with softmax or sigmoid output and to optimize log loss (an observation with predicted probability $p$ incurs a cost of $\log p$, the expectation of which is cross entropy) on the training set with some regularization.
In relational datasets with no negative statements, the prediction that optimizes log loss is to predict everything is true, with a loss of zero on any data set that only includes positive statements. However for a test set that includes statements that should be false, this prediction has an infinite test loss. In training, the lack of negative statements is generally handled by a form of contrastive learning: adding random tuples and using these as negative examples (the few that might be correct are noise). The number of negative versus positive examples provides input that is just fiction. Unlike a prior probability, this input is not overwhelmed by observed data.
Relational models differ from standard tabular models in two main ways:
  • parameter sharing or weight tying; namely “logical variables” and “quantification” in logic approaches (Poole 1993; De Raedt, Kimmig, and Toivonen 2007; Domingos and Lowd 2009), exchangeability [^3] Identifiers can be exchanged and the probability distribution should not change. Exchangeability of identifiers does not imply exchangeability of random variables (de Finetti 1974). (Niepert and Van den Broeck 2014), or the “convolutional” in convolutional graph neural networks (Schlichtkrull et al. 2018).
  • aggregation, where a property of an entity or relation among entities depends on other entities, such as predicting the gender or age of a person from the movies they have watched.
There are a myriad of, often ingenious, learning techniques used for relational data which I won’t go into.
Evaluation
[…]
One could evaluate predictions using log loss, as in the training outlined above. Those papers that report log loss typically do the following: Given a test triple, mutate it by replacing the subject or object to create a negative example. The predictor is then evaluated on these positive and negative examples. (The examples generated to be negative that are actually true are just noise.) Unfortunately the probabilities are just fiction, reflecting the statistics of the knowledge base, not the statistics of the world.
[…]
Ranking has many problems, including: following.
  • Some questions cannot be asked such as “who is the Pope married to?” (where the answer is “nobody”).
  • Asking $(s,v,?)$ leaks information about the test set. Asking who $e$ is married to tells us that $e$ is a person, an adult and not the Pope. Knowing that someone rated a movie, as can be obtained from asking for a prediction for the rating of a movie, is valuable information that can be exploited to produce better predictions. One could imagine a task where the test set is meant to provide information about an entity, as a prompt to an LLM might, and asks questions about that entity. However for many tasks, questions should not provide information. Whether leaking information about the test set is appropriate depends on the downstream application.
  • Some queries are trivial; for example, with fewer than 10 football (soccer) positions and most teams having all of them, asking for a position in a team (see Figure 3), any list of all positions will give a hit-at-10.
  • Some queries are almost impossible, for example asking which soccer team has someone playing “forward” – one of the queries of the FB15k test set – means listing all soccer teams. Guessing which one the query is about is almost impossible. You would expect that an omniscient agent – one who knows everything – would go well in any reasonable evaluation. However, an omniscient agent who knows all soccer teams and all players and their positions, would not be able to guess what team the query was about.
  • The actual probabilities are lost. One predictor might be very sure of one answer, and have to fill in the rest. Another predictor might have a handful of plausible answers, they are not sure about. These predictors could produce the same ordering. There is no penalty for being overconfident or reward for being sure (and correct) or for actually specifying uncertainty, even though this is valuable information for making decisions.
[…]
Looking Forward
There are many success stories from instances of relational learning from protein prediction (Jumper et al. 2021), where the atoms are entities and relations includes bonds, to predicting congestion on roads and at intersections for route-planning (Delling et al. 2015) to the combination of ML and the internet of things in agriculture (Sharma and Shivandu 2024; Mureithi 2024). These build models particular to the application. Generalizing particular models so we can make predictions from arbitrary relational datasets and queries is still in its infancy, despite years of research. This provides a great opportunity for researchers to make fundamental contributions. For relational learning to reach its full potential, the following (at least) need to be done:
We need to consider real public datasets on problems that someone cares about. Unfortunately someone caring means that the datasets are often valuable and so propriety. Relational learning is ideal for making predictions on electronic health records (Weiss et al. 2012; Natarajan et al. 2013), but such records are confidential and cannot be made public for other researchers. Many of the current public datasets are public because no one (currently) cares about them. One class of datasets that should be exploited more is environmental datasets produced and published by government agencies, such as European Environmental Agency [^5] See e.g., https://www.eea.europa.eu/en/datahub. These often contain all measurements taken before a certain time, even if they are not complete with respect to related causes of these. These datasets, however, are often overwhelming for researchers who want to concentrate on the general problem of relational learning, and are much more difficult to use than the datasets in the ML repositories (but see Tschalzev et al. 2025). What is useful to actually predict from these is not obvious for non-domain experts. We need to interact with diverse sets of domain experts to find useful tasks.
The main use of predictions is make decisions. One of the greatest inventions of the 20th century was utility theory (Neumann and Morgenstern 1953): utilities are measures of preference that can be combined with probabilities. To make decisions, we thus need probabilistic predictions and utility functions. For relational models, we need better evaluation to handle the predicament that we want to predict probabilities but only have positive statements. One solution is to include and exploit meta-information, such “there are no more triples involving this subject and verb”. There is a continuum between closed-world and open-world assumptions; datasets are often complete for some narrow part of the world.
For incomplete knowledge bases, we need to explicitly consider why data is missing. Having test examples that are a random subset of all of the tuples, does not provide a surrogate to real-world problems where data is invariably not missing at random (Rubin 1976). I doubt there are any Taylor Swift officially released albums missing from Wikidata, but for most of the recording artists on Wikidata most of their recordings are missing. Probabilistic models of missing data (Marlin et al. 2011; Mohan, Pearl, and Tian 2013) tend to be heavyweight, requiring detailed modelling of why data is missing. This is difficult when virtually all of the tuples are missing. We made an attempt to build a lightweight model of missing data (Poole, Mehr, and Wang 2020) with the aim of extended it to relational models.
When predicting entities, rather than assuming that the correct answer is one of the known entities, there are three possible types of answers: one or more of the known entities, an entity or entities not represented, or that there is no entity. Consider asking for birth mother. For most people in a knowledge base the answer is “someone not represented”. Asking for a child of a person, a person may have (perhaps multiple) children represented, children not represented (or both) or have no children. Similarly, asking for an event that caused something, it might be a known event, an unknown event or there might no event. An unknown event may need to be given an identifier to give it properties. A prediction should not be judged on the actual identier assigned.
We need to be clear about whether we want to predict general knowledge or learn about the entities in the training set. Unless the test sets are carefully designed, there will tend to be a bias towards learn about just one of these. Many models such as probabilistic logic programs (Poole 1993; De Raedt, Kimmig, and Toivonen 2007) and convolutional graph neural networks (Schlichtkrull et al. 2018) learn what looks like general knowledge that is then be applied to the training set to predict properties and relations of the entities in training set. Tests sets from the same dataset do not test whether these models generalize to new populations or related domains. To have statistics, we need to combine the data in at least one dimension; e.g., we could combine temporal data to learn about road congestion at a particular intersection, but could also generalize over the type of intersection. Map builders build models that generalize over only some of the dimensions, but I haven’t seen it for relational learning in general.
Aggregation (e.g., predicting gender of a user from the movies they watched) is the Achilles heel of relational learning. Most models either have a built-in aggregation such as noisy-or for probabilistic logic programs (Poole 1993) or logistic regression for weighted logical formula models (Domingos and Lowd 2009; Kazemi et al. 2014), use a distribution of related entities in relational attention (Velickovic et al. 2018; Wang et al. 2019) or use explicit operations such a sum or mean (Schlichtkrull et al. 2018). They are problematic theoretically (e.g., asymptoticly) (Poole et al. 2014; Buchman and Poole 2015; Adam-Day et al. 2024) and in practice (Kazemi et al. 2017). They need to work with no related entities and as the number of related entities goes to infinity. The models either implicitly assume the related entities provide independent evidence (as do sum, noisy-or and logistic regression), or act the same as if there was just one related entity (as do max, average and attention). Determining whether the evidence is independent (so can be accumulated) or is dependent (in which case extra evidence can be discounted) is difficult, and invariably ignored.
Many methods rely on embeddings/latent features, often fixed-sized embeddings. It seems bizarre to have the same size embeddings for the USA and the relationship between Sinclair and the Portland Ferns (Figure 2). Somehow we need the embedding/latent complexity to depend on the complexity of (the information about) the entity.
To realize the full potential of relational learning, we need to combine the information from all datasets, and build predictive models from these. Combining multiple heterogenous datasets is the challenge of the semantic web (Berners-Lee, Hendler, and Lassila 2001). Building predictive models from heterogenous data might seem impossible, but we know how to; it’s called science. To combine all recorded experimental and observational data to make better predictions, we need at least ontologies [^6] Scientific ontologies include Open Biological and Biomedical Ontology http://obofoundry.org, WHO Family of International Classifications https://www.who.int/standards/classifications, SMOMED-CT healthcare ontology https://www.snomed.org (to allow semantic interoperability) and provenance information (Gil et al. 2016; Sikos, Seneviratne, and McGuinness 2021), including what was manipulated, what was controlled for, what was observed and when, how was what to record decided (modelling the missing information), who manipulated the data, etc. Science involved building hypotheses, typically from multiple datasets, that persist, ready to be revised or rejected as new evidence comes in. Hypotheses can be about anything; from the effects of global warming, to what foods are edible, to the traffic in Vancouver. There needs to be multiple hypotheses for a dataset, including the null hypothesis (the data is noise) and hypotheses that some data is erroneous or entirely fictitious (to counter misinformation). Multiple hypotheses may need to be combined to make a prediction for an actual case (Poole 2012). With an interconnected world, we don’t need to make conclusions from individual experiments, but can learn from everyones experiences, which should enable more accurate hypotheses.
Maybe relational learning in its full generality is impossible, and building separate applications for each domain is the best we do. The only way to determine this is to try to build models that go across domains.
Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it is unfair to count those. Instead we approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I chose the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
For each entity (including reified entities), Wikidata allows the user to download the triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. We define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, we counted the number of local entities that are the subject of fewer than 10 triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 836 for the Sun and 506 for the composer and conductor Pierre Boulez. The lowest of these are just stubs; there are only 11 and 8 tuples respectively with the entity as the subject of a tuple (the 11 included the name and labels in both English and Portuguese). The average number of local entities was 129 (or 108 if the top and bottom outliers were removed). If, for each entity, there are 108 entities that are the subject of fewer than 10 triples, then these entities are 99.08% of the entities.
Wikidata ID # tuples #<10\#<10# < 10 name
Q107132670 67 34 Beverly White
Q3142353 110 50 Hughie Ferguson
Q931229 87 38 USS Congress
Q11813 1055 471 James Madison
Q3830799 157 60 Leroy Chollet
Q6529597 110 49 Les Holden
Q133116210 11 5 History of infant schools in Great Britain
Q4767002 208 89 Anna Filosofova
Q485172 679 283 Ann Arbor
Q16267674 397 277 Hotline Miami 2: Wrong Number
Q1349580 198 92 Matthew Brettingham
Q632511 302 103 Seattle Sounders FC
Q1137583 237 118 Pulgasari
Q3518834 58 20 Territorial Force
Q98076985 29 21 Flight Pattern
Q822900 211 36 geography of Ireland
Q192920 463 160 Edward the Martyr
Q116245633 19 17 Steele’s Greenville expedition
Q525 2588 835 Sun
Q122330832 66 47 All-American Bitch
Q130337417 32 17 The True Record
Q3856283 69 29 Michael Tritter
Q556464 268 156 Gertie the Dinosaur
Q5461994 177 110 Flotilla
Q156193 845 506 Pierre Boulez
Q545151 140 51 The Spy Who Loved Me
Q830590 62 32 Interstate 182
Q3142456 75 41 Hurricane Cindy
Q131151772 9 6 Your Girl
Average 301 129
Figure 4: For each entity, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, and the English name of the entity
Thus for each top-level entity there about 100 other entities with very few tuples, and is of the right order-of-magnitude. There are also top-level entities with few triples. Thus, 99% seems like a reasonable estimate.
As the number of triples increases for an entity, so does the number of entities with few triples. Thus, as a knowledge graph gets bigger and more filled out, the proportion of entities with few triples increases. This goes against the idea that we just need more data.
Note that Wikidata did not include entities for all of the topics of the Wikipedia features articles. The March 12 wikipedia article was for the 2020 season for Seattle Sounders FC, which was not an entity in Wikidata, so we used Seattle Sounders FC (Q632511). There was no Wikidata entry for the March 14 page on the five-pound British gold coin, or the March 31 entry for Apollo 15 postal covers incident. These were omitted. There were 2 entities for “geography of Ireland”, one for the country and one for the island; the featured article was for the island, so that is the entity used.
Appendix B Appendix B: An MRR agent
The mean reciprocal rank (MRR) weights the $n$th result as $\frac{1}{n}$. As in discounting for MDPs and reinforcement learning, this can be interpreted as the expected value for an agent with a probabilistic stopping rule. An MRR agent always considers the first item, considers the second item with probability $\frac{1}{2}$, the third item with probability $\frac{1}{3}$, etc. The MRR of an item corresponds to the probability that such an agent considers the item.
The probability that an MRR agent considers exactly $n$ items is
$\frac{1}{n}-\frac{1}{n+1}=\frac{1}{n(n+1)}.$
The expected number of items this agent considers is
$\sum_{n}n\frac{1}{n(n+1)}=\sum_{n}\frac{1}{n+1}=\infty$
(This is the harmonic series, which is well known to not converge). The problem is that the trillionth (and greater) element needs to be reached with a probability such that computing these large items dominates the time. With finite datasets, the divergence is just of theoretical interest. The agent needs to sometimes consider all of the elements, but sometimes stops just short of this. It is possible that such an agent models a real agent, but I have never seen any evidence that this model fits the behavior of any real agents.
Note that an agent that considers the $n$th item with probability $1/n^{k}$ for $k>1$ does converge (similar to the Riemann zeta function). But then $k$ is an arbitrary parameter which the MRR supposedly avoids.
Companies that present lists of results, such as Google, will know the distribution of the number of items considered. It would be better to use this distribution for evaluation than MRR, but it probably depends on what is being presented (and whether predictions can be tested to determine if that are correct or what is wanted).